Express 15cm as meters using decimals.

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
Meter is a measurement of length. There are 100 centimeters in a meter. 1 meter=100cm, using this convert 15 cm into meters and convert the result into decimals. Complete step-by-step solution: We are given with a length of 15cm and we have to convert it to meters using decimals. We know that 1 meter has 100 centimeters in it. 2 meters has 200 centimeters in it. In the same way, 1m=100cm→1cm=1100m 1 centimeter is one hundredth of one meter. We need to find how many meters is 15cm. 1cm=1100m15cm=?15cm=15×1100m=15100m In decimals, 1100=0.01 Convert 15 cm into decimal meters. 1100=0.0115100=15×1100=15×0.01=0.15 Therefore, 15100m is equal to 0.15m in decimals. 15cm is 0.15 meters in decimal meters. Additional information: A kilometer (km), meter (m), and centimeter (cm) are the commonly used units of length. Any of these measurements can be converted to other measurements using formulas. Additionally, in the customary system (followed in the United States) inches, feet, yards, and miles are used as the unit for length. We can measure how long things are, or how tall, or how far apart they are using the measurements of length. Note: Decimal number can be defined as a number whose whole number part and the fractional part is separated by a decimal point. The dot in a decimal number is called a decimal point. The digits following the decimal point show a value smaller than one. If the whole number part is 0 in a decimal, then the value of the decimal is always less than 1.